# A kiloparsec-scale ordered magnetic field in a galaxy at z=5.6

### v1

@@
Magnetic fields are prevalent on almost all astronomical scales, but their importance in different systems and over cosmic time is yet to be understood. Our current knowledge on the evolution of magnetic fields is limited by scarce observations in the distant Universe, where galaxies have recently been found to be more evolved than most of our model predictions. In this study, we conduct rest-frame 131 m full-polarisation ALMA observations of dust emission in a strongly lensed dusty star-forming galaxy, SPT034652, at , when the Universe was only 1 Gyr old. Dust grains can become aligned with local magnetic fields, resulting in the emission of linearly polarised thermal infrared radiation. Our observations have revealed a median polarisation level of per cent with a variation of per cent across the regions with polarisation detection, similar to that of local starburst galaxies. The polarised dust emission is patchy and mostly overlaps with the [C ii] emission at a velocity of about 150 km s-1, and extends over 3 kiloparsecs. The inferred magnetic field orientations show a bimodal distribution. Our analysis indicates that the kpc-scale polarised dust is most likely aligned by the large-scale ordered magnetic fields associated with a galaxy merger. The early detection of large-scale ordered magnetic fields may favour an efficient formation of magnetic fields in primordial galaxies, which highlights the importance of magnetic fields in mediating galaxy evolution over long cosmic timescales. Future surveys towards a wider galaxy population are necessary to test the ubiquitousness of large-scale magnetic fields in early galaxies.
@@
Magnetic fields are ubiquitous in the multi-phase interstellar medium (ISM) of galaxies (Beck & Wielebinski 2013; Planck Collaboration et al. 2015; Han 2017; Borlaff et al. 2023), but their origin remains a mystery. The development of magnetic fields is generally thought to start from seed fields and then be amplified through various dynamos (see the recent reviews by Subramanian 2019; Brandenburg & Ntormousi 2023). The seed fields could be primordial, either a relic of the early Universe (during the inflation and phase transition) or generated along with the earliest cosmic structure formation, with a strength G (Durrer & Neronov 2013; Subramanian 2019). Beyond this, along with the formation of proto-galaxies, Population-III stars, supernovae, and early active galactic nuclei (AGN) could also deposit up to G seed fields into the ISM (e.g. Hanayama et al. 2005; Xu et al. 2008; Beck et al. 2013; Attia et al. 2021). These seed fields can then be amplified through the small-scale dynamo (also called fluctuation dynamo, turbulent dynamo, Brandenburg & Subramanian 2005), which converts the turbulent kinetic energy of the gas into magnetic field energy until the medium at a given scale saturates (i.e., reach equipartition). In turn, the amplified turbulent fields from the small-scale dynamo undergo an ordering process to form large-scale magnetic fields spanning parsec (pc) to kiloparsec (kpc) scales. Theoretically, the amplification and ordering timescale are sensitive to the model assumptions, which make the observation of early magnetic fields crucial to discriminating different models (e.g. Arshakian et al. 2009; Rodrigues et al. 2019).
@@
Observing magnetic fields back in cosmic time is an ongoing effort. The current most successful method is Faraday rotation; it measures the rotation of the polarisation angle when a background linearly polarised light passes through a magneto-ionic medium. By statistically measuring the Faraday rotation in a sample of high-redshift polarised quasars with and without intervening galaxies, G regular field have been reported in these intervening galaxies up to redshifts of about 2 (Bernet et al. 2008, 2013; Kronberg et al. 2008; Farnes et al. 2014, 2017). Detecting the magnetic fields within individual distant galaxies is more challenging with Faraday rotation because it is difficult to separate the contribution from different components along the line of sight. Recently, Mao et al. (2017) circumvented this difficulty by measuring the differential Faraday rotation between two sub-images of a lensed galaxy, whose light passes through different parts of a foreground late-type galaxy at . Assuming the differential rotation comes mainly from the different field strengths within the foreground lensing galaxy, the Faraday depth indicates G-level coherent magnetic fields around the foreground galaxy. With improved sensitivity of the radio telescope, a similar method has been applied to a larger sample of galaxies with different impact parameters (e.g. Böckmann et al. 2023) and also the host galaxies of fast radio bursts (FRBs) (e.g. Mannings et al. 2023). Although those results are promising, the method can hardly be extended to earlier cosmic time without waiting for the next-generation radio telescopes, e.g., the Square Kilometre Array (SKA) and the next-generation Very Large Array (ngVLA) (Beck et al. 2019; Heald et al. 2020).
@@
Polarised dust is another promising magnetic field tracer across cosmic time. Dust grains are not perfectly spherical; when they are exposed to anisotropic radiation fields, the radiation absorbed by the surface of the dust will create net radiative torque (RAT) to rotate the grains along the minor axis. Dust grains are also paramagnetic, which causes the rotating dust to generate an internal magnetic moment that ends up aligned with the orientation of the local magnetic field. The final configuration is that the dust grain’s minor axis aligns with the local magnetic fields by Larmor precession to minimise the magnetic moment of inertia (Lazarian & Hoang 2007; Hoang & Lazarian 2008, 2016). This mechanism is known as B-RAT, which is the most accepted mechanism driving the grain alignment in the ISM (Andersson et al. 2015; Lopez-Rodriguez & Tram 2024). In nearby galaxies, far-infrared (FIR) polarimetric observations have been the most efficient way to probe the geometry of the projected magnetic fields in the cold phase of the interstellar medium (Lopez-Rodriguez et al. 2021, 2022b, 2022a; Pattle et al. 2021; Lopez-Rodriguez 2023) and it has also been used to measure the strength of galactic magnetic fields in starburst galaxies with large-scale ordered galactic flows (Lopez-Rodriguez et al. 2021) and in gas streamers in the centre of our Galaxy (Guerra et al. 2023).
@@
Dust is ubiquitous among star-forming galaxies (SFGs). The advent of submillimetre/millimetre (submm/mm) telescopes has pushed the detections of dust emission back to the epoch of cosmic reionisation, benefiting from the negative- correction (e.g. Laporte et al. 2017; Tamura et al. 2019; Inami et al. 2022). Meanwhile, hundreds of the most bright dusty SFGs have been confirmed to be strongly lensed systems, with their apparent flux densities and spatial scales being amplified by a foreground galaxy or galaxy cluster (e.g. Reuter et al. 2020; Urquhart et al. 2022). The lens amplification can facilitate the detection of faint polarised dust emission and boost the spatial resolution. In addition, polarised thermal dust emission is not affected by Faraday rotation. All these advantages make dust polarisation a promising option to map the galactic magnetic fields across cosmic time with modern submm/mm interferometers. The current highest-redshift detection of polarised dust emission in a lensed galaxy, 9io9, at has been recently reported by Geach et al. (2023). The dust polarisation fraction in 9io9 is of the order of 1 per cent and extends over 5 kpc. Such a similar polarisation fraction compared to nearby star-forming galaxies is intriguing, considering that 9io9 has an order of magnitude higher gas mass and star formation rate (SFR). The ordered magnetic fields are generally aligned with the cold gas disk, favouring a rapid formation of ordered magnetic fields. Such a success also opens a new window to explore magnetic fields in distant galaxies.
@@
In this paper, we conduct full-polarisation ALMA dust continuum observation of SPT0346-52 at an even higher redshift, z5.56 (1 Gyr after the Big Bang, the Cosmic Dawn), to map its galactic magnetic fields with polarised dust emission. SPT034652 is the brightest and most intensely star-forming galaxy discovered in the 2,500 deg2 survey of the South Pole Telescope (Vieira et al. 2010; Hezaveh et al. 2013). Lens modelling shows that SPT034656 has been magnified by a foreground galaxy (at ) with an amplification factor of and a half-light radius of kpc (Spilker et al. 2016). Correcting for the lens amplification, SPT034652 is an intrinsic hyper-luminous infrared galaxy (HyLIRG, L) and has a star-formation rate (SFR) around (Ma et al. 2016). Given its compact size, SPT034652 has one of the most extreme SFR surface densities (Ma et al. 2016). In addition, SPT034652 has a highly enriched ISM, dominated by warm and dense molecular gas (Apostolovski et al. 2019; Litke et al. 2022). The total flux density of SPT034652 at m is mJy (Reuter et al. 2020), which makes it the best target to detect and resolve the polarised dust emissions back to the Cosmic Dawn.
@@
The paper is organised as follows. In section 2, we describe the basic information about the full-polarisation observations and data reduction. In section 3, we derive the linear polarisation properties of the dust emission and the geometry of the inferred magnetic fields. In section 4, we compare the observed results with the nearby galaxies and discuss possible physical origins of the ordered magnetic fields, their formation timescale, and their implications for the early galaxy evolution. Finally, we summarise our work in section 5. A flat CDM cosmology with , , and km s-1 Mpc-1 is assumed throughout this work.
@@ 2 Observation and data reduction
The data were calibrated with CASA (ver.6.2.1, The CASA Team et al. 2022). We first followed the standard calibration pipeline of the two parallel polarisations, XX and YY, wherein J05194546 and J02535441 were used to calibrate the bandpass and phase variation, respectively. The flux of all the targets was calibrated from the flux calibrator J05194546, with an uncertainty of about 5 per cent. Then, the calibrated data were split for polarisation calibration. We followed the official template to apply the polarisation calibration. For both sessions, the blazar J04230120 was used as the polarisation calibrator to correct the instrumental polarisation leakage. The best-fit polarisation fraction of J04230120 is around 4.0 per cent, consistent with the long-term monitoring of its polarisation status 222https://www.alma.cl/ skameno/AMAPOLA/J0423-0120.flux.html. The instrumental uncertainty of the linear polarised fraction is about 0.1 per cent and 1 deg in terms of polarisation angle from ALMA technical handbook (Cortes et al. 2023).
After the polarisation calibration, we used task tclean within CASA to make the images and cubes. We first created image cubes of the Stokes I (the total intensity) to identify possible spectral lines. The cubes were created with the ‘hogbom’ deconvolver and natural weighting. We did not identify any strong lines () within our spectral windows. Then, we created the full Stokes images with the ‘clarkstokes’ deconvolver and natural weighting to maximise the signal-to-noise-ratio (S/N). All the clean processes stop when the residuals reach the root mean squared (RMS) level of 12 Jy beam-1 with a beam size of ( kpc at =5.6). The RMS achieved for the polarised intensity is 8.5 Jy beam-1.
@@ 3 Results and Analysis
The inferred magnetic fields with Nyquist sampling are shown in Figs. 1 and 2, overlaid on the total and polarised intensity map, respectively. The histograms of the polarisation fraction, , and magnetic field position angle, , are also shown in Fig. 2. The vertical dashed and dotted lines indicate the median and mean of polarisation fractions among all the samplings. The median and mean values of polarisation fractions both show per cent, with 0.4 per cent variations across the individual polarisation measurements. The measured polarisation fraction is close to the measurement in 9io9, where a mean fraction of per cent is reported at rest-frame 350 m and extends over 5 kpc parallel to the galactic disk (Geach et al. 2023). The position angles of the magnetic fields in SPT034652 have a large variation, ranging from 0 to 100 degrees, and show a bimodal distribution. The vertical dash-dotted lines show the two peaks of the bimodal distribution at 28 degrees and 78 degrees, with median polarisation fractions of and per cent, respectively.
To recover the dust emission in the source plane, we used a pixellated lensing reconstruction technique developed by Warren & Dye (2003) and used extensively in the field with the public available software (PyAutoLens, Nightingale & Dye 2015; Nightingale et al. 2021; Dye et al. 2018, 2022). This method makes no assumption about the source structure, allowing complex source features to be recovered. To better guide the lens modelling, we mask the regions without signals, including the central area with 0.05 arcsec and the outer area with 0.18 arcsec. From our best-fit model, the Stokes of SPT034652 is reconstructed to a single elongated and relatively compact region of emission in the source plane, which is consistent with the existing lens models (e.g. Spilker et al. 2015; Litke et al. 2019; Jones et al. 2019a). However, due to the faintness of the polarised dust emission, we can only marginally recover its structure in the source plane (see more details in Fig. 3). The recovered field structures favour the presence of ordered magnetic fields in the outskirts of the lensed galaxy. The details about full-polarisation lens modelling and comparing different techniques, including the modelling in the visibility space, will be covered in the forthcoming paper by Liu et al. (in prep.). The reconstructed source models are shown in Fig. 3 and the best-fit lens parameters are summarised in Tab. 2. The physical resolution of our observation varies at different galactic positions. The region with the polarisation detection has the highest amplification, with a mean amplification factor of , which corresponds to pc in the source plane with our angular resolution. After correcting the lens amplification, the whole region extends over 3 kpc in the source plane.

### v2

@@
Magnetic fields are prevalent on almost all astronomical scales, but their importance in different systems and over cosmic time is yet to be understood. Our current knowledge of the evolution of magnetic fields is limited by scarce observations in the distant Universe, where galaxies have recently been found to be more evolved than most model predictions. In this study, we conducted rest-frame 131 m full-polarisation observations of dust emission in a strongly lensed dusty star-forming galaxy, SPT034652, at , when the Universe was only 1 Gyr old. Dust grains can become aligned with local magnetic fields, resulting in the emission of linearly polarised thermal infrared radiation. Our observations have revealed a median polarisation level of per cent with a variation of per cent across the regions with polarisation detection, similar to that of local starburst galaxies. The polarised dust emission is patchy and mostly overlaps with the [C ii] emission at a velocity of about 150 km s-1, and extends over 3 kiloparsecs. The bimodal distribution of field orientations, their spatial distribution, and the connection with the cold gas kinematics further emphasise the complexity of the magnetic environment in this galaxy and the potential role of mergers in shaping its magnetic fields. Such early formation of ordered galactic magnetic fields also suggests that both small-scale and large-scale dynamos could be efficient in early galaxies. Continued observations of magnetic fields in early galaxies, as well as expanding surveys to a wider galaxy population, are essential for a comprehensive understanding of the prevalence and impact of magnetic fields in the evolving Universe.
@@
Magnetic fields are ubiquitous in the multi-phase interstellar medium (ISM), and circum-galactic medium (CGM) of galaxies (Beck & Wielebinski 2013; Planck Collaboration et al. 2015; Han 2017; Borlaff et al. 2023), but their origin remains a mystery. The development of magnetic fields is generally thought to start from seed fields and then be amplified through various dynamos (see the recent reviews by Subramanian 2019; Brandenburg & Ntormousi 2023). The seed fields could be primordial, either a relic of the early Universe (during the inflation and phase transition) or generated along with the earliest cosmic structure formation, with a strength G (Durrer & Neronov 2013; Subramanian 2019). Beyond this, along with the formation of proto-galaxies, Population-III stars, supernovae, and early active galactic nuclei (AGN) could also deposit up to G seed fields into the ISM (e.g. Hanayama et al. 2005; Xu et al. 2008; Beck et al. 2013; Attia et al. 2021). These seed fields can then be amplified through the small-scale dynamo (also called fluctuation dynamo, turbulent dynamo, Brandenburg & Subramanian 2005), which converts the turbulent kinetic energy of the gas into magnetic field energy until the medium at a given scale saturates (i.e. reach equipartition). In turn, the amplified turbulent fields from the small-scale dynamo undergo ordering processes, such as - dynamo (Ruzmaikin et al. 1988; Beck 2015; Brandenburg & Ntormousi 2023), to form large-scale magnetic fields spanning sub-kiloparsec (kpc) to tens of kpc scales. Theoretically, the amplification and ordering timescale are sensitive to the model assumptions, which make the observation of early magnetic fields crucial to discriminating different models (e.g. Arshakian et al. 2009; Rodrigues et al. 2019).
@@
The current most explored method is Faraday rotation; it measures the rotation of the polarisation angle when a background linearly polarised light passes through a magneto-ionic medium. By statistically measuring the Faraday rotation in a sample of high-redshift polarised quasars with and without intervening galaxies, G regular field (parallel to the line-of-sight) have been reported in these intervening galaxies up to redshifts of about 2 (e.g. Bernet et al. 2008, 2013; Kronberg et al. 2008; Farnes et al. 2014, 2017). However, the total rotation measures integrate all the contributions along the line of sight, including the ISM/CGM of the quasars, intervening material, and Galactic magnetic fields. Different samples or different models of the Galactic magnetic fields can sometimes lead to controversial results (Oren & Wolfe 1995; Kim et al. 2016; Lan & Prochaska 2020; Malik et al. 2020). Recently, Mao et al. (2017) circumvented this difficulty by measuring the differential Faraday rotation between two sub-images of a lensed quasar, whose light passes through different parts of a foreground late-type galaxy at . Assuming the differential rotation comes mainly from the different field strengths within the foreground lensing galaxy, the Faraday depth indicates G-level regular magnetic fields around the foreground galaxy. With improved sensitivity of the radio telescopes, the Faraday rotation measurements have been applied to an increasingly larger number of intervening galaxies towards the sightlines of the quasars with different impact parameters (e.g. Böckmann et al. 2023). Meanwhile, in addition to distant quasars, the fast radio bursts (FRBs) have become another promising background polarised source (e.g. Mannings et al. 2023). All these progresses show a growing community of mapping magnetic fields in galaxies on various spatial scales and at different cosmic epochs. More importantly, the next-generation radio telescopes, e.g. the Square Kilometre Array (SKA) and the next-generation Very Large Array (ngVLA)(Beck et al. 2019; Heald et al. 2020), can further test these results and push the measurements to an earlier cosmic time.
@@
Polarised dust is another promising magnetic field tracer across cosmic time. Dust grains are not perfectly spherical; when they are exposed to anisotropic radiation fields, the radiation absorbed by the surface of the dust will create net radiative torque (RAT) to rotate the grains along the minor axis. Dust grains are also paramagnetic, which causes the rotating dust to generate an internal magnetic moment that ends up aligned with the orientation of the local magnetic field. The final configuration is that the dust grain’s minor axis aligns with the local magnetic fields by Larmor precession to minimise the magnetic moment of inertia (Lazarian & Hoang 2007; Hoang & Lazarian 2008, 2016). This mechanism is known as B-RAT, which is the most accepted mechanism driving the grain alignment in the ISM (Andersson et al. 2015; Lopez-Rodriguez & Tram 2024). In nearby galaxies, far-infrared (FIR) polarimetric observations have been the most efficient way to probe the geometry of the projected magnetic fields in the cold phase of the interstellar medium (Lopez-Rodriguez et al. 2021, 2022a, 2022b; Pattle et al. 2021; Lopez-Rodriguez 2023). To differentiate the magnetic fields measured with different techniques, hereafter, we will use “ordered” fields to represent the inferred magnetic fields from polarised dust emission, which has 180 degrees of ambiguity, and use “regular” for fields with a well-defined direction.
@@
Dust is ubiquitous among star-forming galaxies (SFGs). The advent of submillimetre/millimetre (submm/mm) telescopes has pushed the detections of dust emission back to the epoch of cosmic reionisation, benefiting from the negative- correction (e.g. Laporte et al. 2017; Tamura et al. 2019; Inami et al. 2022). Meanwhile, hundreds of the most bright dusty SFGs have been confirmed to be strongly lensed systems, with their apparent flux densities and spatial scales being amplified by a foreground galaxy or galaxy cluster (e.g. Reuter et al. 2020; Urquhart et al. 2022). The lens amplification can facilitate the detection of faint polarised dust emission and boost the spatial resolution. In addition, polarised thermal dust emission is not affected by Faraday rotation. All these advantages make dust polarisation a promising option to map the galactic magnetic fields across cosmic time with modern submm/mm interferometers. The current highest-redshift detection of polarised dust emission in a lensed galaxy, 9io9, at has been recently reported by Geach et al. (2023). The dust polarisation fraction in 9io9 is of the order of 1 per cent and the polarised dust emission extends over 5 kpc along the galaxy’s disk. The polarisation fraction in 9io9 is comparable to nearby star-forming galaxies, which is intriguing considering that 9io9 has an order of magnitude higher gas mass and star formation rate (SFR). The ordered magnetic fields are generally aligned with the cold gas disk, favouring a rapid formation of ordered magnetic fields. Such a success also opens a new window to explore magnetic fields in distant galaxies.
@@
In this paper, we conduct full-polarisation Atacama Large Millimeter/Submillimeter Array (ALMA) dust continuum observation of 2 at an even higher redshift, z=5.56 ( Gyr after the Big Bang, the Cosmic Dawn), to map its galactic magnetic fields with polarised dust emission. SPT034652 is the brightest and most intensely star-forming galaxy discovered in the 2,500 deg2 survey of the South Pole Telescope (Vieira et al. 2010; Hezaveh et al. 2013). Lens modelling shows that SPT034656 has been magnified by a foreground galaxy (at ) with an amplification factor of and a half-light radius of kpc (Spilker et al. 2016). Correcting for the lens amplification, SPT034652 is an intrinsic hyper-luminous infrared galaxy (HyLIRG, L) and has a SFR around (Ma et al. 2016). Given its compact size, SPT034652 has one of the most extreme SFR surface densities (Ma et al. 2016). In addition, SPT034652 has a highly enriched ISM, dominated by warm and dense molecular gas (Apostolovski et al. 2019; Litke et al. 2022). The total flux density of SPT034652 at m is mJy (Reuter et al. 2020), which makes it the best target to detect and resolve the polarised dust emissions back to the Cosmic Dawn.
@@
The paper is organised as follows. In section 2, we describe the basic information about the full-polarisation observations and data reduction. In section 3, we derive the linear polarisation properties of the dust emission and the geometry of the inferred magnetic fields. In section 4, we compare the observed results with the nearby galaxies and discuss possible physical origins of the ordered magnetic fields, their formation timescale, and their implications for early galaxy evolution. Finally, we summarise our work in section LABEL:sec:conclusion. A flat CDM cosmology with , , and km s-1 Mpc-1 is assumed throughout this work.
@@ 2 Observation and data reduction
The data were calibrated with Common Astronomy Software Applications (CASA) package (ver.6.2.1, The CASA Team et al. 2022). We first followed the standard calibration pipeline of the two parallel polarisations, XX and YY, wherein J05194546 and J02535441 were used to calibrate the bandpass and phase variation, respectively. The flux of all the targets was calibrated from the flux calibrator J05194546, with an uncertainty of about 5 per cent. Then, the calibrated data were split for polarisation calibration. We followed the official template to apply the polarisation calibration. For both sessions, the blazar J04230120 was used as the polarisation calibrator to correct the instrumental polarisation leakage. The best-fit polarisation fraction of J04230120 is around 4.70.1 per cent, consistent with the long-term monitoring of its polarisation status of 3.2-5.1 per cent at the same period 222https://www.alma.cl/ skameno/AMAPOLA/J0423-0120.flux.html. The measured uncertainty in polarisation angle is 0.2 deg. The instrumental uncertainty of the linear polarised fraction is about 0.1 per cent and 1 deg in terms of polarisation angle from ALMA technical handbook (Cortes et al. 2023).
After the polarisation calibration, we used task tclean within CASA to make the images and cubes. We first created image cubes of the Stokes I (the total intensity) to identify possible spectral lines. The cubes were created with the ‘hogbom’ deconvolver and natural weighting. We did not identify any strong lines () within our spectral windows. Then, we created the full Stokes images with the ‘clarkstokes’ deconvolver and natural weighting to maximise the signal-to-noise ratio (S/N). All the clean processes stop when the residuals reach the root mean squared (RMS) level of 12 Jy beam-1 with a beam size of ( kpc at =5.6). The RMS achieved for the polarised intensity is 8.5 Jy beam-1.
@@ 3 Results and analysis
The inferred magnetic fields with Nyquist sampling are shown in Figs. 1 and 2, overlaid on the total and polarised intensity map, respectively. The histograms of the polarisation fraction, , and magnetic field position angle, , are also shown in Fig. 2. The vertical dashed and dotted lines indicate the median and mean of polarisation fractions among all the samplings. The median and mean values of polarisation fractions both show per cent, with 0.4 per cent variations across the individual polarisation measurements. The measured polarisation fraction is close to the measurement in 9io9, where a mean fraction of per cent is reported at rest-frame 350 m and extends over 5 kpc parallel to the galactic disk (Geach et al. 2023). The position angles of the magnetic fields in SPT034652 have a large variation, ranging from degrees, and show a bimodal distribution. The vertical dash-dotted lines show the two peaks of the bimodal distribution at 28 degrees and 78 degrees, with median polarisation fractions of and per cent, respectively.
To recover the dust emission in the source plane, we used a pixellated lensing reconstruction technique developed by Warren & Dye (2003) and used extensively in the field with the public available software (PyAutoLens, Nightingale & Dye 2015; Nightingale et al. 2021; Dye et al. 2018, 2022). This method makes no assumption about the source structure, allowing complex source features to be recovered. To better guide the lens modelling, we masked the central area with 0.05 arcsec and the outer area with 0.18 arcsec. The reconstructed source models are shown in Fig. 3 and the best-fit lens parameters are summarised in Tab. 2. The Stokes of SPT034652 is reconstructed to a single elongated and relatively compact region of emission in the source plane, which is consistent with the existing lens models (e.g. Spilker et al. 2015; Litke et al. 2019; Jones et al. 2019a). The recovered spatial structure of the polarised emission is offset roughly 300-600 pc from the peak of Stokes and only present in the north and east sides of the lensed galaxy. The position angles of the field line change gradually from the north side to the east side, which favours the presence of ordered magnetic fields in these regions. However, due to the faintness of the polarised dust emission, we can only marginally recover its structure in the source plane by modelling the region with possible signals. We are exploring different ways to model lensed full-polarisation data, including the modelling in the visibility space. The details will be reported in the forthcoming paper by Liu et al. (in prep.). The physical resolution of our observation varies at different galactic positions. The region with the polarisation detection has the highest amplification, with a mean amplification factor of , which corresponds to pc in the source plane with our angular resolution. After correcting the lens amplification, the polarised dust extends over 3 kpc in the source plane.